# Temperature-Dependent Emission Spectroscopy of Quantum Emitters in Hexagonal Boron Nitride

### v1

@@ I Introduction
Color centers in hBN Tran et al. (2016a) have emerged as an excellent source of single photons for applications in quantum technologies Ahmadi et al. (2024); Çakan et al. (2025); Mai et al. (2025); Cholsuk et al. (2024a). Due to the intrinsically large bandgap of approximately 6 eV in hBN Cassabois et al. (2016), various color centers can be formed and, in turn, create additional energy levels in the bandgap. Depending on the transition energies between the two-level defect states of each color center, these single photon emitters (SPEs) can potentially generate single photons across a broad spectral range Bourrellier et al. (2016); Shevitski et al. (2019); Cholsuk et al. (2022, 2024b); Camphausen et al. (2020). Many of these emitters exhibit long-term stability Vogl et al. (2018), high brightness at room temperature Tran et al. (2016a), narrow linewidths Dietrich et al. (2018), and weak electron-phonon coupling Li et al. (2017), making them suitable for integration into quantum photonic devices. In addition, the atomically thin, layered structure of hBN facilitates efficient photon collection Vogl et al. (2019) and integration into on-chip platforms Nonahal et al. (2023); Li et al. (2021). Despite these advantages, preselection of the properties of hBN emitters, such as their emission wavelength, polarization direction, and spatial localization remains challenging. The microscopic nature of many hBN defect centers is to-date poorly understood Cholsuk et al. (2023). While the same fabrication techniques can produce multiple species of emitters Hayee et al. (2020); Tran et al. (2016b), the emission properties of a single emitter can vary due to external factors such as temperature Jungwirth et al. (2016); Korkut and Sarpkaya (2022), excitation energy Schell et al. (2018); Shotan et al. (2016), or local strain in the crystal lattice Grosso et al. (2017). Therefore, harnessing the full potential of hBN SPEs requires further elucidation of their electronic and optical properties.
@@ II.2 Photophysical Properties at Cryogenic Temperature
After identifying the defect-related emission, we investigated its photophysical properties at low temperature. The emitter was continuously excited non-resonantly with an 80 MHz pulsed laser. The resulting emission spectra over a 50-minute period are presented in Fig. 2(a). The ZPL exhibited spectral diffusion around 547.5 $\pm$ 0.2 nm as depicted in the inset. Importantly we note that no lines appear/ disappear, which would reveal additional energy levels or more complicated photodynamics. Meanwhile, time-resolved photon counting over 5 minutes revealed no signs of blinking or bleaching, as can be seen in Fig. 2(b,c). The mean count rate is 2290 $\pm$ 172 counts/s, which remains above the shot noise limit (the square-root of the mean is 48 counts/s).
@@ II.3 Temperature-Dependent Photoluminescence of the Emitter
To quantitatively analyze the temperature-induced changes in the ZPL, the emission peak was fitted using a Lorentzian function (see Fig. A3), and both the full-width at half-maximum (FWHM) and the central wavelength were extracted. The results of the fitting are presented in Fig. 3(b) and (c) respectively. In the low-temperature regime (up to approximately 50 K), the ZPL linewidth remains nearly constant, suggesting that dephasing is dominated by temperature-independent mechanisms. Beyond that, the ZPL linewidth increases with temperature, as depicted in Fig. 3(b). Similarly, the ZPL emission energy remains nearly unchanged below 50 K and begins to redshift by approximately 6 meV between 4.5 and 200 K. This shift can be attributed to reduced bandgap at elevated temperature O’donnell and Chen (1991); Du et al. (2014); Tan et al. (2022). To understand the underlying phonon contributions, the temperature dependence of the FWHM and peak energy of ZPL was fitted to various power-law functions. In the intermediate temperature range (approximately 50–150 K), the FWHM follows a $T^{3}$ trend that has been observed previously in hBN Jungwirth et al. (2016); Tan et al. (2022); Korkut and Sarpkaya (2022); White et al. (2021); Sontheimer et al. (2017) and other color centers Neu et al. (2013); Mueller et al. (2012); Lienhard et al. (2016); Hizhnyakov et al. (2002). Meanwhile, the ZPL redshift above 50 K is well described by a $T^{2}$ dependence. These temperature dependencies deviate from the predictions of standard harmonic models Maradudin (1966). Specifically, quadratic electron-phonon coupling under the Debye approximation anticipates a $T^{7}$ dependence for linewidth broadening and a $T^{4}$ dependence for ZPL redshift. Furthermore, an increase in linewidth at higher temperatures slower than $T^{5}$ indicates that dynamic Jahn-Teller effects are unlikely to play a significant role Fu et al. (2009). Taken together, these deviations suggest alternative dephasing mechanisms. Notably, the observed $T^{3}$ linewidth broadening and $T^{2}$ redshift are consistent with soft-mode theory Hizhnyakov et al. (2002); Mueller et al. (2012); Razgulov et al. (2021); Neu et al. (2013). In this framework, electronic excitation weakens the local bonding, increasing the density of low-frequency phonon modes Hizhnyakov et al. (2002). As temperature rises, the Bose–Einstein occupation of these phonon modes increases, enhancing their coupling to the electronic states. This leads to temperature-dependent dephasing and energy renormalization, observed as ZPL broadening and redshift.
@@ III Conclusions
In summary, we investigated low-temperature PL from exfoliated hBN samples irradiated with electron beam using an SEM. By systematically varying the excitation power and photon energy of the laser, we identified defect-related emission at cryogenic temperature. Furthermore, the use of tunable excitation allowed us to distinguish between Raman scattering and defect-related luminescence, enabling more accurate attribution of spectral features. Following the identification of the defect-related emission, we characterized the photophysical properties of the emitters. The emission remained stable over time, showing minimal blinking and spectral diffusion under cryogenic conditions. Additionally, we investigated the temperature-dependent PL of electron-irradiated hBN to gain insights into the electron-phonon interactions governing the emission properties. Similar studies have previously been conducted on hBN quantum emitters, revealing various temperature-dependent behaviors of ZPL linewidth, such as $T^{3}$ White et al. (2021); Sontheimer et al. (2017); Tan et al. (2022), $T+T^{5}$ Arı et al. (2025), nearly exponential Jungwirth et al. (2016), or linear trends Akbari et al. (2021) implying different underlying physical mechanisms. The emitters in those studies were typically created/activated via annealing drop-casted hBN nanoflakes in ethanol/water solution in N2 environment Sontheimer et al. (2017); Tan et al. (2022); Jungwirth et al. (2016), by exposing exfoliated hBN to H2 White et al. (2021) plasma followed by annealing in air and annealing in an Ar environment Akbari et al. (2021). Differences observed across the literature are likely attributable to the wide range of structural and electronic configurations of defect centers, highlighting the intrinsic heterogeneity of color centers in hBN. While low-temperature studies on hBN emitters created via electron beam irradiation have been reported Gale et al. (2022); Fournier et al. (2021); Zhigulin et al. (2023b), our result represents a comprehensive temperature-dependent investigation of their optical properties. Deeper insight into these systems is essential for the development of quantum technologies, where characteristics such as emission linewidth, spectral stability, and photon indistinguishability are important. In addition, identification of the phonon mode, together with PLE measurements, revealed an efficient excitation mechanism via optical phonon-assisted absorption, leading to a nearly fivefold enhancement in emission. Despite the presence of strong electron-phonon coupling, which tends to suppress the ZPL at elevated temperatures, the emitters demonstrated high single-photon purity even at room temperature, as indicated by low $g^{2}(0)$ values in our previous work Kumar et al. (2023, 2024a). This makes them promising candidates for quantum key distribution systems Abasifard et al. (2024), where both emission probability and photon purity are critical.
@@ Acknowledgements
This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111-390814868 (MCQST) and as part of the CRC 1375 NOA project C2 and project PQET (INST 95/1654-1). The authors acknowledge support from the Federal Ministry of Education and Research (BMBF) under grant number 13N16292 (ATOMIQS) and by the German Space Agency DLR with funds provided by the Federal Ministry for Economic Affairs and Climate Action BMWK under grant numbers 50WM2165 (QUICK3), and 50RP2200 (QuVeKS).

### v2

@@ I Introduction
Color centers in hBN Tran et al. (2016a) have emerged as an excellent source of single photons for applications in quantum technologies Ahmadi et al. (2024); Çakan et al. (2025); Mai et al. (2025); Cholsuk et al. (2024a). Due to the intrinsically large bandgap of approximately 6 eV in hBN Cassabois et al. (2016), various color centers can be formed and, in turn, create additional energy levels in the bandgap. Depending on the transition energies between the two-level defect states of each color center, these single photon emitters (SPEs) can potentially generate single photons across a broad spectral range Bourrellier et al. (2016); Shevitski et al. (2019); Cholsuk et al. (2022, 2024b); Camphausen et al. (2020). Many of these emitters exhibit long-term stability Vogl et al. (2018), high brightness at room temperature Tran et al. (2016a), narrow linewidths Dietrich et al. (2018), and weak electron-phonon coupling Li et al. (2017), making them suitable for integration into quantum photonic devices. In addition, the atomically thin, layered structure of hBN facilitates efficient photon collection Vogl et al. (2019) and integration into on-chip platforms Nonahal et al. (2023); Li et al. (2021). Despite these advantages, preselection of the properties of hBN emitters, such as their emission wavelength, polarization direction, and spatial localization remains challenging. The microscopic nature of many hBN defect centers is to-date poorly understood Cholsuk et al. (2023). While the same fabrication techniques can produce multiple species of emitters Hayee et al. (2020); Tran et al. (2016b), the emission properties of a single emitter can vary due to external factors such as temperature jungwirth2016temperature; Korkut and Sarpkaya (2022), excitation energy Schell et al. (2018); Shotan et al. (2016), or local strain in the crystal lattice Grosso et al. (2017). Therefore, harnessing the full potential of hBN SPEs requires further elucidation of their electronic and optical properties.
@@ II.2 Photophysical Properties at Cryogenic Temperature
After identifying the defect-related emission, we investigated its photophysical properties at low temperature. The emitter was continuously excited non-resonantly with an 80 MHz pulsed laser. The resulting emission spectra over a 50-minute period are presented in Fig. 2(a). The ZPL exhibited spectral diffusion around 547.5 $\pm$ 0.2 nm as depicted in the inset. Integrating the spectra for all time steps and fitting it in a Gaussian yields a linewidth of 260 GHz. Importantly we note that no lines appear/ disappear, which would reveal additional energy levels or more complicated photodynamics. Meanwhile, time-resolved photon counting over 5 minutes revealed no signs of blinking or bleaching, as can be seen in Fig. 2(b,c). The mean count rate is 2290 $\pm$ 172 counts/s, which remains above the shot noise limit (the square-root of the mean is 48 counts/s).
@@ II.3 Temperature-Dependent Photoluminescence of the Emitter
To quantitatively analyze the temperature-induced changes in the ZPL, the emission peak was fitted using a Lorentzian function (see Fig. A3), and both the full-width at half-maximum (FWHM) and the central wavelength were extracted. The results of the fitting are presented in Fig. 3(b) and (c) respectively. In the low-temperature regime (up to approximately 50 K), the ZPL linewidth remains nearly constant, suggesting that dephasing is dominated by temperature-independent mechanisms. Beyond that, the ZPL linewidth increases with temperature, as depicted in Fig. 3(b). Similarly, the ZPL emission energy remains nearly unchanged below 50 K and begins to redshift by approximately 6 meV between 4.5 and 200 K. This shift can be attributed to reduced bandgap at elevated temperature O’donnell and Chen (1991); Du et al. (2014); Tan et al. (2022). To understand the underlying phonon contributions, the temperature dependence of the FWHM and peak energy of ZPL was fitted to various power-law functions. In the intermediate temperature range (approximately 50–150 K), the FWHM follows a $T^{3}$ trend that has been observed previously in hBN jungwirth2016temperature; Tan et al. (2022); Korkut and Sarpkaya (2022); White et al. (2021); Sontheimer et al. (2017) and other color centers Neu et al. (2013); Mueller et al. (2012); Lienhard et al. (2016); Hizhnyakov et al. (2002). Meanwhile, the ZPL redshift above 50 K is well described by a $T^{2}$ dependence. These temperature dependencies deviate from the predictions of standard harmonic models Maradudin (1966). Specifically, quadratic electron-phonon coupling under the Debye approximation anticipates a $T^{7}$ dependence for linewidth broadening and a $T^{4}$ dependence for ZPL redshift. Furthermore, an increase in linewidth at higher temperatures slower than $T^{5}$ indicates that dynamic Jahn-Teller effects are unlikely to play a significant role Fu et al. (2009). Taken together, these deviations suggest alternative dephasing mechanisms. Notably, the observed $T^{3}$ linewidth broadening and $T^{2}$ redshift are consistent with soft-mode theory Hizhnyakov et al. (2002); Mueller et al. (2012); Razgulov et al. (2021); Neu et al. (2013). In this framework, electronic excitation weakens the local bonding, increasing the density of low-frequency phonon modes Hizhnyakov et al. (2002). As temperature rises, the Bose–Einstein occupation of these phonon modes increases, enhancing their coupling to the electronic states. This leads to temperature-dependent dephasing and energy renormalization, observed as ZPL broadening and redshift.
@@ III Conclusions
In summary, we investigated low-temperature PL from exfoliated hBN samples irradiated with electron beam using an SEM. By systematically varying the excitation power and photon energy of the laser, we identified defect-related emission at cryogenic temperature. Furthermore, the use of tunable excitation allowed us to distinguish between Raman scattering and defect-related luminescence, enabling more accurate attribution of spectral features. Following the identification of the defect-related emission, we characterized the photophysical properties of the emitters. The emission remained stable over time, showing minimal blinking and spectral diffusion under cryogenic conditions. Additionally, we investigated the temperature-dependent PL of electron-irradiated hBN to gain insights into the electron-phonon interactions governing the emission properties. Similar studies have previously been conducted on hBN quantum emitters, revealing various temperature-dependent behaviors of ZPL linewidth, such as $T^{3}$ White et al. (2021); Sontheimer et al. (2017); Tan et al. (2022), $T+T^{5}$ Arı et al. (2025), nearly exponential jungwirth2016temperature, or linear trends Akbari et al. (2021) implying different underlying physical mechanisms. The emitters in those studies were typically created/activated via annealing drop-casted hBN nanoflakes in ethanol/water solution in N2 environment Sontheimer et al. (2017); Tan et al. (2022); jungwirth2016temperature, by exposing exfoliated hBN to H2 White et al. (2021) plasma followed by annealing in air and annealing in an Ar environment Akbari et al. (2021). Differences observed across the literature are likely attributable to the wide range of structural and electronic configurations of defect centers, highlighting the intrinsic heterogeneity of color centers in hBN. While low-temperature studies on hBN emitters created via electron beam irradiation have been reported Gale et al. (2022); Fournier et al. (2021); Zhigulin et al. (2023b), our result represents a comprehensive temperature-dependent investigation of their optical properties. Deeper insight into these systems is essential for the development of quantum technologies, where characteristics such as emission linewidth, spectral stability, and photon indistinguishability are important. In addition, identification of the phonon mode, together with PLE measurements, revealed an efficient excitation mechanism via optical phonon-assisted absorption, leading to a nearly fivefold enhancement in emission. Despite the presence of strong electron-phonon coupling, which tends to suppress the ZPL at elevated temperatures, the emitters demonstrated high single-photon purity even at room temperature, as indicated by low $g^{2}(0)$ values in our previous work Kumar et al. (2023, 2024a). This makes them promising candidates for quantum key distribution systems Abasifard et al. (2024), where both emission probability and photon purity are critical.
@@ Acknowledgements
This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111-390814868 (MCQST) and as part of the CRC 1375 NOA project C2 and project PQET (INST 95/1654-1). The authors acknowledge support from the Federal Ministry of Research, Technology and Space (BMFTR) under grant number 13N16292 (ATOMIQS) and by the German Space Agency DLR with funds provided by the Federal Ministry for Economic Affairs and Climate Action BMWK under grant numbers 50WM2165 (QUICK3), and 50RP2200 (QuVeKS).